# Improved QLDPC Surgery: Logical Measurements and Bridging Codes

### v3

@@ 3.3 Code Distance and Fault Distance
Though we proved it in the special case of measuring $\bar{X}_{M}$ for simplicity, this lemma can be generalized to apply to any of the ancilla measurement schemes in this paper. There is always an original stabilizer group $\mathcal{S}_{O}$ and merged code group $\mathcal{S}_{M}$ measured $R\geq 1$ times. Let $\mathcal{U}=\langle\mathcal{S}_{O},\mathcal{S}_{M}\rangle$. If $\bar{P}_{M}$ is the logical operator being measured, let $\bar{Q}_{M}$ be a Pauli that anticommutes with it. We can establish a group of logical operators for $k-1$ qubits $\mathcal{L}$ that is the same for all the codes, and such that elements of $\mathcal{L}$ commute with both $\bar{P}_{M}$ and $\bar{Q}_{M}$. The lemma then applies with $\bar{Q}_{M}$ replacing $\bar{Z}_{M}$. For instance, if we are measuring $\bar{P}_{M}=\bar{X}_{1}\bar{X}_{2}$ as in the next section, we can take $\bar{Q}_{M}=\bar{Z}_{1}$, set up $\mathcal{L}=\langle\bar{X}_{2},\bar{Z}_{1}\bar{Z}_{2},\bar{X}_{i},\bar{Z}_{i},\forall i=3,...k\rangle$, and calculate $d(\bar{Q}_{M}\mathcal{L}\mathcal{U})$ and $d(\mathcal{L}^{*}\mathcal{U})$ to bound the measurement and logical fault distances.
@@
We note that initializing the module code on qubits $V_{i\geq 2\text{ even}}\cup C_{j\geq 3\text{ odd}}$ was used in our $\bar{X}_{M}$ measurement protocol just to make a specific choice of codes at the merge and split steps. This is flexible to some extent. For instance, one could instead initialize all qubits $V_{i\geq 2\text{ even}}\cup C_{j\geq 1\text{ odd}}$ to |$\ket{0}$ and make the same arguments to arrive at Lemma 9 and Theorem 11. The key property we need from a choice of initialization is that the $Z$ checks are reliable at the merge and split.
@@ 3.4.1 Hypergraph Product Codes
$X_{n\times n}$ =$=\{(w\otimes r,0_{m\times m}):w\in W\subset\ker(H),r\in R\subset\mathbb{F}_{2}^{n}/\operatorname{im}(H^{\top})\},$ (28)
$X_{m\times m}$ =$=\{(0_{n\times n},v\otimes s):v\in V\subset\ker(H^{\top}),s\in S\subset\mathbb{F}_{2}^{m}/\operatorname{im}(H)\}.$ (29)
We cite as a fact that these two sets specify a complete basis of the $X$-logical operators of $Q$. Note that $|X_{n\times n}|=(n-r)^{2}$ and $|X_{m\times m}|=(m-r)^{2}$.
Theorem 13.
For a gauge-fixed ancilla system on an $X$-operator $\bar{X}\in X_{n\times n}\cup X_{m\times m}$, all gauge checks are redundant.
Proof.
Fix an $X$-operator $P$ supported on $(w\otimes r,0_{m\times m})\in X_{n\times n}$, we will prove this claim for the gauge-fixed ancilla system built on $P$. The case where $P$ is supported on a basis vector in $X_{m\times m}$ follows similarly.
@@
Let $C=\{f_{j}\in\mathbb{F}_{2}^{m}:H^{\top}f_{j}\cap w\neq\varnothing\}$, then $C\otimes r=\{f\otimes r:f\in C\}$ is exactly the set of $Z$-checks of $Q$ that touches $V_{0}$. Let $F$ be the submatrix of $H$ induced by $C$ and $\operatorname{supp}(w)$, namely, $F:C\rightarrow\operatorname{supp}(w)$. Since $r$ is a standard basis vector, we have
$F=F\otimes r:C\otimes r\rightarrow\operatorname{supp}(w\otimes r)$ (31)
is the submatrix of $H_{Z}$ induced by the qubits $V_{0}$ and $Z$-checks $C_{0}=C\otimes r$. Therefore, in a $L$ layer gauge-fixed ancilla system build on $P$, we can identify every $C_{i}$ as a copy of $C$, and every $V_{i}$ as a copy $\operatorname{supp}(w)$.
The gauge $Z$-check of our system corresponds to vectors $c\in\text{null}(F)\subset C$, i.e., $cF=0$. Fix a gauge check $Z(c)$, and consider its equivalent representative supported on qubits in $C_{1}$. We will show that $Z(c)$ is a product of $Z$-stabilizers in $\mathcal{C}_{Z}$. Let $\mathcal{H}_{Z}$ denote the new $Z$-check matrix with the ancilla system. Consider the product of $Z$-checks $c\otimes r\in\mathcal{C}^{Z}$, which is
@@
Since $c\in C$ is in null space of $F$, and $F$ is the submatrix of $H$ induced by $C$ and $\operatorname{supp}(w)$, we must have $H^{\top}c=0$. Therefore,
@@
Note that the first equation is true because the $Z$-checks in $c\otimes v$ are kept unchanged by the ancilla system, which is only attached to the checks $C\otimes r$. This proves that the gauge check $Z(c\in C_{1})$ is a product of $Z$-stabilizers in $\mathcal{C}^{Z}$. ∎
As a corollary, for hypergraph product codes the merged code $\mathcal{G_{X}}$ has degree $\delta+1$ if the original code has degree $\delta$. We expect that similar analysis can be made for lifted product codes and potentially other families of QLDPC codes. We further remark that Lemma 12 extends to all the joint measurement systems which we discuss later.
@@ 3.5 Modular Decoder Achieving Measurement Fault-Distance
Above we showed that the phenomenological fault-distances of the decoding graph $\mathcal{D}$ are $d$ for both measurement errors and logical errors whenever it is constructed with at least $R\geq d$ rounds. To ensure practicality of the scheme we must also ensure that a decoder can correct errors of weight $<d/2$. It is reasonable to assume that such a decoder already exists for the code $\mathcal{G}$ on its own. In this section we show that if there is a good decoder for the original code $\mathcal{G}$, then there also exists a good decoder for extracting measurement outcomes from the measurement protocol specified by $\mathcal{D}$.
@@
Suppose there exists a decoder that, for any collection $E_{\mathcal{G}}$ of $Z$-qubit and $X$-measurement errors of weight $<d/2$ in the detector graph $\mathcal{D}^{\mathcal{G}}$, outputs a correction $E^{\prime}_{\mathcal{G}}$ consisting of $Z$-qubit errors and $X$-measurement errors consistent with the detectors and equivalent to $E_{\mathcal{G}}$ up to the stabilizers of $\mathcal{G}$.
Then there exists a decoder for $\mathcal{D}$ as a whole that outputs a correction $E^{\prime}$ such that, if the true error is $E$ of weight $<d/2$, the product $EE^{\prime}$ triggers no detectors and commutes with the measured observable $\bar{X}_{M}$.
@@
We select the matching that minimizes the length between paired steps. For example, the lengths of the matchings when $k$ is odd are $(j_{2}-j_{1})+(j_{4}-j_{3})+(R-j_{k})$, and $(j_{1}-0)+(j_{2}-j_{1})+(j_{k}-j_{k-1})$ respectively.
@@
Suppose the true error on $\mathcal{D}$ was $E$ of weight $<d/2$. It remains to show that $EE^{\prime}$ commutes with the measured observable $\bar{X}$. $\bar{X}$ is the product of all the $X$-checks on $V_{j}$ for odd $j$. Here we point out that several parts of $E$ introduced above do not affect the measurement result. Hence a worst-case error $E$ does not feature these parts, since they needlessy increase the weight. This lets us restrict our attention to corrections $E$ of a certain form. Since $E_{\mathcal{G}}E^{\prime}_{\mathcal{G}}$ is equivalent to the identity up to the stabilizer of $\mathcal{G}$ by assumption, we can restrict our attention to an error $E$ of weight $<d/2$ supported on $\mathcal{D}^{\text{anc}}$ only. Furthermore, since the errors identified from Lemma 8 flip an even number of $X$-checks in any $V_{j}$ for odd $j$, these must commute with $\bar{X}$ as well. So we may furthermore assume that $E$ flips at most one detector for each $V_{j}$ within each layer. Similarly, flipping an even number of bits within the odd $V_{j}$ groups within an even layer also commutes with the measurement outcome, so can assume that errors only occur on $V_{1}$.
@@
We observe that if every qubit in $C_{1}$ is adjacent to exactly two checks from $V_{1}$, then we can simply use a minimum-weight perfect matching (MWPM) decoder on $\mathcal{D}^{\text{anc}}$. This condition is equivalent to every $Z$-check in $C_{0}$ touching two qubits in $V_{0}$. This is not true for arbitrary codes and logical operators. However, in our case study on the $\llbracket 144,12,12\rrbracket$ code in Section 4, we find logical operators where this condition is satisfied.
@@ 3.6 Joint $X$ Measurement
Our method of joint measurement is distinct from the proposal in [CKBB22], not just from the fact that our ancilla systems are gauge-fixed, but also because their approach actually requires $12d-$ 1 layers to preserve the code distance and, more specifically, to preserve the weight of $\bar{X}_{1}$ or $\bar{X}_{2}$, which are still logical operators in the merged code. Instead of more layers, we introduce a collection of $\min(|\bar{X}_{1}|,|\bar{X}_{2}|)$ qubits that are connected to $X$ checks in both ancilla systems. We term these qubits bridge qubits. The joint system is depicted in Figure 4.
@@
A complete, linearly independent basis of gauge checks of $U^{B}$ in the joint ancilla system consists of $|B|-1$ checks. Moreover, there is a basis in which each check acts on exactly two bridge qubits.
@@
With the help of these gauge checks, we can show that the merged code indeed has $k-1$ logical qubits.
@@
The merged code $\mathcal{G}_{XX}$ for measuring the product of two logical $X$ operators has $k-1$ logical qubits.
@@
First, notice that the original code $\mathcal{G}$ with ancilla systems for measuring $\bar{X}_{1}$ and $\bar{X}_{2}$ would encode $k-2$ logical qubits by applying Theorem 1 twice. Since $\mathcal{G}_{XX}$ adds $|B|$ qubits and $|B|-1$ independent checks (see Lemma 15) to that code, it encodes $k-1$ logical qubits. ∎
In practice, the joint measurement can be carried out with a schedule similar to the $X$ measurement, Section 3.2 – starting from the original code, measure the stabilizers of the joint merged code for $R$ rounds, and finally measure the stabilizers of the original code again. The qubits in $V_{i}$ for even $i\geq 2$, $C_{j}$ for $j\geq 1$ odd, and in $B$ can be initialized in |$\ket{0}$ at the zeroth round and measured out in the $Z$ basis at the final round. Note this dispenses with the idea of preparing the module in a stabilizer state before the protocol and preserving it in that state after the protocol in favor of simpler single-qubit preparations and final measurements. However, we can reach the same conclusions below if the modules for measuring $\bar{X}_{1}$ and $\bar{X}_{2}$ were kept intact as well.
@@
Let $\mathcal{S}$ be the stabilizer group of the original code and each ancilla qubit, including bridge qubits, in |$\ket{0}$, and $\mathcal{S}_{XX}$ be the stabilizer group of the joint merged code, Figure 4. Also let $\mathcal{L}$ be the group of logical operators of the unmeasured $k-1$ qubit logical space and =$\mathcal{L}^{*}=\mathcal{L}\setminus\{I\}$. The versions of Lemmas 9 and 10 that apply to the joint code imply that the measurement fault distance is $\min(R,d_{Z}(\bar{Z}_{1}\mathcal{L}\mathcal{S}))$ and the logical fault distance is $\min(d_{Z}(\mathcal{L}^{*}\mathcal{S}),d_{X}(\mathcal{L}^{*}\mathcal{S}_{XX}))$. The original code $\mathcal{S}$ has code distance $d$, therefore $d_{Z}(\bar{Z}_{1}\mathcal{L}\mathcal{S})\geq d$ and $d_{Z}(\mathcal{L}^{*}\mathcal{S})\geq d$. To complete the proof, it remains only to show that the $X$-distance of the joint merged code is good, i.e. $d_{X}(\mathcal{L}^{*}\mathcal{S}_{XX})\geq d$.
@@ 3.7 Gauge-fixed $Y$ Ancilla System
Given the $X$ and $Z$ logical measurement schemes, in this section we present a $Y$ measurement scheme that combines the two. Our proposal differs from the one presented in [CKBB22] because that proposal relies on using $12d-$ 1 layers to achieve fault-distance $d$.
@@
It is important to note that if $\bar{X}_{M}$ and $\bar{Z}_{M}$ overlap on more than one qubit, the above proof may no longer hold, and the $Y$ merged code may have extra gauge qubits. These gauge checks take the form of $X(u\in C_{1}^{X})Z(v\in C_{1}^{Z})$, where $u$ and $v$ are chosen to ensure commutation with checks in $V_{1}^{X}$, $V_{1}^{Z}$, and the new collection of checks $Q_{1}$, $|Q_{1}|>1$, which replaces $q_{1}$. In particular, $Z(v\in C_{1}^{Z})$ flips an even number of checks in $V_{1}^{X}\cup Q_{1}$ and $X(u\in C_{1}^{X})$ flips an even number of checks in $V_{1}^{Z}\cup Q_{1}$. If the only checks they flip are both the same set of checks from $Q_{1}$, then $X(u\in C_{1}^{X})Z(v\in C_{1}^{Z})$ is a gauge operator. This also explains why there are no gauge checks of this form if $Q_{1}=\{q_{1}\}$, because it is impossible to flip an even number of checks from a set of size one. In general, these gauge checks should be fixed as well to ensure the system encodes just $k-1$ qubits.
Knowing the structure of the gauge checks allows us to prove that the $Y$ system encodes exactly $k-1$ qubits, though we defer the proof to Appendix A.2.
@@
The $Y$-system encodes $k-1$ qubits.
To show that the $Y$-system has good fault-distance, we apply Lemma 9 with $\mathcal{U}=\langle\mathcal{S}_{O},\mathcal{S}_{Y}\rangle$ and $\mathcal{L}=\langle\bar{X}_{i},\bar{Z}_{i},i=1,2,\dots,k-1\rangle$, where $\mathcal{S}_{O}$ is the stabilizer group of the original code (with the additional qubits starting in $\ket{+}$ for the bottom half of Figure 5 and the bridge and in |$\ket{0}$ for the top half), $\mathcal{S}_{Y}$ the stabilizer group of the merged code shown in Figure 5, and $\mathcal{L}$ a simultaneous $k-1$ qubit logical basis for each of the codes $\mathcal{S}_{O}^{\prime}=\langle\mathcal{S}_{O},\bar{Y}_{M}\rangle$, $\mathcal{S}_{Y}$ and $\mathcal{U}$. We have that the logical fault distance is at least $d(\mathcal{L}^{*}\mathcal{U})$ and the measurement fault distance is $\min(R,d(\bar{Z}_{M}\mathcal{L}\mathcal{U}),d(\bar{X}_{M}\mathcal{L}\mathcal{U}))$ for $R$ rounds of measurement.
@@
$\operatorname{Init}=\langle Z(v)\text{ for singleton }v\in C^{Z}_{j\text{ odd}},X(v)\text{ for singleton }v\in C^{X}_{j\text{ odd}},X(b)\text{ for singleton }b\in B\rangle,$ (61)
where we used $C^{Z}_{j\text{ odd}}$ to mean $\bigcup_{j\text{ odd}}C^{Z}_{j}$ and similarly for the other variants of this notation. Then we have

### v4

@@ 3.3 Code Distance and Fault Distance
Though we proved it in the special case of measuring $\bar{X}_{M}$ for simplicity, this lemma can be generalized to apply to any of the ancilla measurement schemes in this paper. There is always an original stabilizer group $\mathcal{S}_{O}$ and merged code group $\mathcal{S}_{M}$ measured $R\geq 1$ times. Let $\mathcal{U}=\langle\mathcal{S}_{O},\mathcal{S}_{M}\rangle$. If $\bar{P}_{M}$ is the logical operator being measured, let $\bar{Q}_{M}$ be a Pauli that anticommutes with it. We can establish a group of logical operators for k1k-1 qubits $\mathcal{L}$ that is the same for all the codes, and such that elements of $\mathcal{L}$ commute with both $\bar{P}_{M}$ and $\bar{Q}_{M}$. The lemma then applies with $\bar{Q}_{M}$ replacing $\bar{Z}_{M}$. For instance, if we are measuring $\bar{P}_{M}=\bar{X}_{1}\bar{X}_{2}$ as in the next section, we can take $\bar{Q}_{M}=\bar{Z}_{1}$, set up $\mathcal{L}=\langle\bar{X}_{2},\bar{Z}_{1}\bar{Z}_{2},\bar{X}_{i},\bar{Z}_{i},\forall i=3,...k\rangle$, and calculate $d(\bar{Q}_{M}\mathcal{L}\mathcal{U})$ and $d(\mathcal{L}^{*}\mathcal{U})$ to bound the measurement and logical fault distances.
@@
[…]
@@ 3.4 Redundancy of Gauge Checks
$X_{n\times n}$ $=\{(w\otimes r,0_{m\times m}):w\in W\subset\ker(H),r\in R\subset\mathbb{F}_{2}^{n}/\operatorname{im}(H^{\top})\},$ (28)
$X_{m\times m}$ $=\{(0_{n\times n},s\otimes v):s\in S\subset\mathbb{F}_{2}^{m}/\operatorname{im}(H),v\in V\subset\ker(H^{\top})\}.$ (29)
We cite as a fact that these two sets specify a complete basis of the $X$-logical operators of $Q$. Note that $|X_{n\times n}|=(n-r)^{2}$ and $|X_{m\times m}|=(m-r)^{2}$. Fix an $X$-operator $P$ supported on $(w\otimes r,0_{m\times m})\in X_{n\times n}$, we will analyze the gauge-fixed ancilla system built on $P$. The case where $P$ is supported on a basis vector in $X_{m\times m}$ follows similarly.
@@
Let $C=\{f_{j}\in\mathbb{F}_{2}^{m}:H^{\top}f_{j}\cap w\neq\varnothing\}$, then $C\otimes r=\{f\otimes r:f\in C\}$ is exactly the set of $Z$-checks of $Q$ that touches $V_{0}$. Let $F$ be the submatrix of $H$ induced by $C$ and $\operatorname{supp}(w)$, namely, $F:C\rightarrow\operatorname{supp}(w)$. Since $r$ is a standard basis vector, we see that $F$ is also the submatrix of $H_{Z}$ induced by the qubits $V_{0}$ and $Z$-checks $C_{0}=C\otimes r$. Therefore, in a $L$ layer gauge-fixed ancilla system build on $P$, we can identify every $C_{i}$ as a copy of $C$, and every $V_{i}$ as a copy of $\operatorname{supp}(w)$.
The gauge $Z$-check of our system corresponds to vectors $c\in\text{null}(F)\subset C$, i.e., $cF=0$. We consider a subset of these gauge checks which corresponds to vectors in $\text{null}(H)$. Specifically, for every $u\in\text{null}(H)$, consider the restriction of $u$ to the indices in $C$, which we denote by $u|_{C}$. By construction of $C$, we see that $u|_{C}F=0$ which means $u|_{C}\in\text{null}(F)$. We show that the gauge checks corresponding to these vectors are redundant.
Fix a vector $c\in C$ such that $c=u|_{C}$ for some $u\in\text{null}(H)$. Consider the gauge check $Z(c)$ and its equivalent representative supported on qubits in $C_{1}$. Let $\mathcal{H}_{Z}$ denote the new $Z$-check matrix with the ancilla system. Consider the product of $Z$-checks $c\otimes r\in\mathcal{C}^{Z}$, which is
@@
Since $H^{\top}c=0$, we have
@@
Note that the first equation is true because the $Z$-checks in $c\otimes v$ are kept unchanged by the ancilla system, which is only attached to the checks $C\otimes r$. This proves that the gauge check $Z(c\in C_{1})$ is a product of $Z$-stabilizers in $\mathcal{C}^{Z}$, and is therefore redundant. We expect that similar analysis can be made for lifted product codes and potentially other families of QLDPC codes.
In an earlier version of this paper, we mistakenly claimed that all gauge checks $Z(c)$ for $c\in\text{null}(F)$ are redundant. We thank John Blue for pointing out an error in the earlier proof, which is now corrected.
@@ 3.5 Modular Decoder Achieving Measurement Fault-Distance
Above we showed that the phenomenological fault-distances of the decoding graph $\mathcal{D}$ are $d$ for both measurement errors and logical errors whenever it is constructed with at least $R\geq d$ rounds. To ensure practicality of the scheme we must also ensure that a decoder can correct errors of weight <d/2<d/2. It is reasonable to assume that such a decoder already exists for the code $\mathcal{G}$ on its own. In this section we show that if there is a good decoder for the original code $\mathcal{G}$, then there also exists a good decoder for extracting measurement outcomes from the measurement protocol specified by $\mathcal{D}$.
@@
Suppose there exists a decoder that, for any collection $E_{\mathcal{G}}$ of $Z$-qubit and $X$-measurement errors of weight <d/2<d/2 in the detector graph $\mathcal{D}^{\mathcal{G}}$, outputs a correction $E^{\prime}_{\mathcal{G}}$ consisting of $Z$-qubit errors and $X$-measurement errors consistent with the detectors and equivalent to $E_{\mathcal{G}}$ up to the stabilizers of $\mathcal{G}$.
Then there exists a decoder for $\mathcal{D}$ as a whole that outputs a correction $E^{\prime}$ such that, if the true error is $E$ of weight <d/2<d/2, the product $EE^{\prime}$ triggers no detectors and commutes with the measured observable $\bar{X}_{M}$.
@@
We select the matching that minimizes the length between paired steps. For example, the lengths of the matchings when $k$ is odd are $(j_{2}-j_{1})+(j_{4}-j_{3})+...+(R-j_{k})$, and $(j_{1}-0)+(j_{2}-j_{1})+...+(j_{k}-j_{k-1})$ respectively.
@@
Suppose the true error on $\mathcal{D}$ was $E$ of weight <d/2<d/2. It remains to show that $EE^{\prime}$ commutes with the measured observable $\bar{X}$. $\bar{X}$ is the product of all the $X$-checks on $V_{j}$ for odd $j$. Here we point out that several parts of $E$ introduced above do not affect the measurement result. Hence a worst-case error $E$ does not feature these parts, since they needlessy increase the weight. This lets us restrict our attention to corrections $E$ of a certain form. Since $E_{\mathcal{G}}E^{\prime}_{\mathcal{G}}$ is equivalent to the identity up to the stabilizer of $\mathcal{G}$ by assumption, we can restrict our attention to an error $E$ of weight <d/2<d/2 supported on $\mathcal{D}^{\text{anc}}$ only. Furthermore, since the errors identified from Lemma 8 flip an even number of $X$-checks in any $V_{j}$ for odd $j$, these must commute with $\bar{X}$ as well. So we may furthermore assume that $E$ flips at most one detector for each $V_{j}$ within each layer. Similarly, flipping an even number of bits within the odd $V_{j}$ groups within an even layer also commutes with the measurement outcome, so can assume that errors only occur on $V_{1}$.
@@
Remark 14.
We note that while the decoding strategy described above correctly achieves decoding distance $\lfloor d/2\rfloor$, as stated it does not have a threshold. The reason is that under stochastic noise, the parity of each step $\mathcal{D}^{\text{anc}}_{i}$ is odd with probability converging to $1/2$, which means the matching decoder on a line will fail with probability converging to $1/2$. This is a shortcoming of the proof, but not of the modular decoding method itself. We can instead use minimum weight decoders on both $\mathcal{D}^{\mathcal{G}}$ and $\mathcal{D}^{\text{anc}}$, which we expect to have a threshold asymptotically.
We further observe that if every qubit in $C_{1}$ is adjacent to exactly two checks from $V_{1}$, then we can simply use a minimum-weight perfect matching (MWPM) decoder on $\mathcal{D}^{\text{anc}}$. This condition is equivalent to every $Z$-check in $C_{0}$ touching two qubits in $V_{0}$. This is not true for arbitrary codes and logical operators. However, in our case study on the $\llbracket 144,12,12\rrbracket$ code in Section 4, we find logical operators where this condition is satisfied. We therefore employ the modular decoder with BPOSD and MWPM in our numerical simulations presented in Section 4.
@@ 3.6 Joint $X$ Measurement
Our method of joint measurement is distinct from the proposal in [CKBB22], not just from the fact that our ancilla systems are gauge-fixed, but also because their approach actually requires 2d12d-1 layers to preserve the code distance and, more specifically, to preserve the weight of $\bar{X}_{1}$ or $\bar{X}_{2}$, which are still logical operators in the merged code. Instead of more layers, we introduce a collection of $\min(|\bar{X}_{1}|,|\bar{X}_{2}|)$ qubits that are connected to $X$ checks in both ancilla systems. We term these qubits bridge qubits. The joint system is depicted in Figure 4.
@@
A complete, linearly independent basis of gauge checks of $U^{B}$ in the joint ancilla system consists of |B|1|B|-1 checks. Moreover, there is a basis in which each check acts on exactly two bridge qubits.
@@
With the help of these gauge checks, we can show that the merged code indeed has k1k-1 logical qubits.
@@
The merged code $\mathcal{G}_{XX}$ for measuring the product of two logical $X$ operators has k1k-1 logical qubits.
@@
First, notice that the original code $\mathcal{G}$ with ancilla systems for measuring $\bar{X}_{1}$ and $\bar{X}_{2}$ would encode k2k-2 logical qubits by applying Theorem 1 twice. Since $\mathcal{G}_{XX}$ adds $|B|$ qubits and |B|1|B|-1 independent checks (see Lemma 15) to that code, it encodes k1k-1 logical qubits. ∎
[…]
@@
[…]
@@ 3.7 Gauge-fixed $Y$ Ancilla System
Given the $X$ and $Z$ logical measurement schemes, in this section we present a $Y$ measurement scheme that combines the two. Our proposal differs from the one presented in [CKBB22] because that proposal relies on using 2d12d-1 layers to achieve fault-distance $d$.
@@
It is important to note that if $\bar{X}_{M}$ and $\bar{Z}_{M}$ overlap on more than one qubit, the above proof may no longer hold, and the $Y$ merged code may have extra gauge qubits. These gauge checks take the form of $X(u\in C_{1}^{X})Z(v\in C_{1}^{Z})$, where $u$ and $v$ are chosen to ensure commutation with checks in $V_{1}^{X}$, $V_{1}^{Z}$, and the new collection of checks $Q_{1}$, $|Q_{1}|>1$, which replaces $q_{1}$. In particular, $Z(v\in C_{1}^{Z})$ flips an even number of checks in $V_{1}^{X}\cup Q_{1}$ and $X(u\in C_{1}^{X})$ flips an even number of checks in $V_{1}^{Z}\cup Q_{1}$. If the only checks they flip are both the same set of checks from $Q_{1}$, then $X(u\in C_{1}^{X})Z(v\in C_{1}^{Z})$ is a gauge operator. This also explains why there are no gauge checks of this form if $Q_{1}=\{q_{1}\}$, because it is impossible to flip an even number of checks from a set of size one. In general, these gauge checks should be fixed as well to ensure the system encodes just k1k-1 qubits.
Knowing the structure of the gauge checks allows us to prove that the $Y$ system encodes exactly k1k-1 qubits, though we defer the proof to Appendix A.2.
@@
The $Y$-system encodes k1k-1 qubits.
[…]
@@
[…]